OpenStudy (anonymous):

cual es la integral de 3x+1/x

OpenStudy (zzr0ck3r):

\[\int\frac{3x+1}{x}dx\]?

OpenStudy (anonymous):

yeah thats what the person is asking for

OpenStudy (zzr0ck3r):

\[\int\frac{3x+1}{x}dx=\int3dx+\int x^{-1}dx=3x+ln(x)+c\]

OpenStudy (anonymous):

gracias por quitarme la duda..